Deshaun invested his savings into two investment funds. The amount he invested in Fund A was $2000 less than the amount he invested in Fund B. find A returned a 5% profit and Fund B returned a 4% profit. How much did he invest in Fund B, if the total profit from the two funds together was $980?

Answer by ikleyn(52787)

.

    0.04*B + 0.05*(B-2000) = 980  dollars.


Simplify and find B

    0.04B + 0.05B - 100 = 980

         0.09B          = 980 + 100 = 1080

             B                      = 1080/0.09 = 12000  dollars.    ANSWER

Solved.

All equations are self-explanatory.